OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random . what is the probability that both numbers are odd? (A) 1/3 (B) 1/2 (C)1/4 (D) 1/8 . can somebody help me

OpenStudy (campbell_st):

P(odd) = 1/2 P(odd, odd) = P(odd) x P(odd)

OpenStudy (raden):

the event of odd numbers are 1,3,5,7,9. so there are 9 events, and there are 10 all events. then use the formula : P(A and B) = P(A) * P(B) P(odd and odd) = 1/2 * 1/2 = 1/4

OpenStudy (anonymous):

yes I believe is 1/4 correct cuz 1/2 I got it rong on my math homework thank you guys